# appIReferences for Extended Technical Appendix

### v3

@@ Abstract
Generating diverse, all-atom conformational ensembles of dynamic proteins such as G-protein-coupled receptors (GPCRs) is critical for understanding their function, yet most generative models simplify atomic detail or ignore conformational diversity altogether. We present latent diffusion for full protein generation (LD-FPG), a framework that constructs complete all-atom protein structures, including every side-chain heavy atom, directly from molecular dynamics (MD) trajectories. LD-FPG employs a Chebyshev graph neural network (ChebNet) to obtain low-dimensional latent embeddings of protein conformations, which are processed using three pooling strategies: blind, sequential and residue-based. A diffusion model trained on these latent representations generates new samples that a decoder, optionally regularized by dihedral-angle losses, maps back to Cartesian coordinates. Using D2R-MD, a $2\,\mu\text{s}$ MD trajectory (12 000 frames) of the human dopamine D2222 receptor in a membrane environment, the sequential and residue-based pooling strategy reproduces the reference ensemble with high structural fidelity (all-atom lDDT $\sim 0.7$; $C\alpha$-lDDT $\sim 0.8$) and recovers backbone and side-chain dihedral-angle distributions with a Jensen–Shannon divergence $<0.03$ compared to the MD data. LD-FPG thereby offers a practical route to system-specific, all-atom ensemble generation for large proteins, providing a promising tool for structure-based therapeutic design on complex, dynamic targets. The D2R-MD dataset and our implementation are freely available to facilitate further research.
@@ 1 Introduction
Proteins function as dynamic molecular machines whose biological activities critically depend on transitioning between distinct conformational states [1, 2]. Landmark artificial intelligence methods like AlphaFold2 [3] and others [4, 5, 6, 7] have advanced structure prediction but predominantly predict single static conformations, limiting their utility for systems with conformational heterogeneity. Accurate modeling of an ensemble of accessible conformations is essential to elucidate protein function and guide therapeutic design [8, 9, 10, 11]. Crucially, these ensembles must explicitly represent all atomic details, particularly side chains, whose subtle conformational rearrangements often govern molecular recognition and catalytic mechanisms [12].
@@
Despite rapid advancements, existing generative models frequently fall short in capturing the detailed dynamics of side-chain movements specific to particular proteins [13]. Many powerful methodologies have focused either on de novo backbone designs [14, 15, 16] or on generating static all-atom structures [17, 18], but neither produce comprehensive conformational ensembles. Ensemble-generating approaches—via perturbations of static predictors [19, 20], flow-matched variants [21], or general MD-trained generators [22, 23]—typically operate at the backbone or coarse-grained level. Consequently, they fail to capture the intricate all-atom rearrangements, particularly involving side chains, that are critical for function. Although promising latent space [24, 25, 26] and physics-informed models [27, 28] have emerged, their capability to generate high-resolution all-atom ensembles from MD data reflecting functional transitions remains unproven, highlighting a significant unmet need for specialized generative frameworks.
@@
G protein-coupled receptors (GPCRs) provide a compelling example of dynamic systems wherein precise all-atom modeling is indispensable [29]. This large family of transmembrane receptors, comprising over 800 human members, is responsible for mediating most known transmembrane signal transduction [30, 31] and is targeted by approximately 50% of all marketed drugs [32, 33]. GPCR signaling involves conformational transitions among multiple states, frequently induced by ligand binding [34], and occurs through intricate allosteric mechanisms where specific side-chain interactions are crucial [35, 36]. Capturing such dynamic, atomically detailed landscapes is vital to understanding receptor signaling specificity [37], biased agonism [38], and the design of drugs targeting unique allosteric sites [39, 40, 41], yet studying these events computationally remains formidable [42, 43, 44]. Current predictive methods for GPCR dynamics [45] do not generate the comprehensive all-atom conformational landscapes essential for mechanistic understanding.
@@
To address this critical requirement, we introduce Latent Diffusion for Full Protein Generation (LD-FPG), a generative framework designed to learn and generate diverse, all-atom conformational ensembles from existing MD simulation data of a target protein, explicitly including side-chain details. Rather than simulating new trajectories, LD-FPG leverages extensive MD datasets. Our approach employs a Chebyshev Spectral Graph Convolutional Network (ChebNet) [46] to encode all-atom MD snapshots into a compact latent representation. A Denoising Diffusion Probabilistic Model (DDPM) [47] is then trained to explore this learned latent manifold, and latent representations are decoded back into full all-atom Cartesian coordinates. We demonstrate our framework on extensive MD simulations of the human dopamine D2 receptor (D2R), systematically evaluating distinct decoder pooling strategies. Our primary contributions in this paper are:
@@ 2 Related Work
Generative Models for Protein Design and Static Structure Prediction. Deep generative models have made significant strides in protein science, employing techniques such as diffusion [47], flow-matching [48, 49], and learned latent spaces [24, 50] to tackle complex structural tasks. For de novo backbone design, various methods, often inspired by the success of predictors like AlphaFold2 [3] and diffusion-based approaches like RFdiffusion [14] have emerged. Notable examples include FoldingDiff [15], alongside flow-matching models such as FoldFlow2 [49]. Latent space strategies have also been pivotal in this domain; for instance, LatentDiff [24] generates novel protein backbones using an equivariant diffusion model within a condensed latent space, while Ig-VAE [50] utilizes a variational autoencoder for class-specific backbone generation (e.g., for immunoglobulins). While powerful for creating new folds (e.g., Proteina [16]) or specific components, these methods typically target soluble proteins, often generate only backbone coordinates, and are not primarily designed for sampling multiple conformations of a specific, existing protein. The generation of complete static all-atom structures has also seen considerable progress. Models like Protpardelle [17] and Chroma [18] can produce full static structures from sequence information. Diffusion-based generative models such as AlphaFold3 [6] and Boltz-1 [7] also provide detailed single-state predictions of all-atom structures and complexes. Other approaches, like PLAID [25], integrate predictors with diffusion samplers. To complement backbone or static generation, methods including FlowPacker [51] and SidechainDiff [52] focus on side-chain packing or prediction. However, these tools predominantly yield single static structures. Furthermore, decoupling backbone and side-chain generation risks overlooking their critical interplay during the complex dynamic transitions [29] relevant to the conformational landscapes our work aims to capture.
@@
Modeling Protein Conformational Diversity: From General Strategies to MD-Informed Approaches. Beyond single static structures, capturing a protein’s conformational diversity is crucial for understanding its function. Initial strategies to explore this diversity include learning from structural variations in experimental databases (e.g., Str2Str [20]) or perturbing static predictions to sample conformational space, as seen with AF2-RAVE [19] for GPCRs and AlphaFlow/ESMFlow [21] more broadly. While these methods effectively broaden sampling, capturing system-specific, native-like dynamics often benefits from models trained directly on simulation data, which can provide a richer representation of a given protein’s accessible states. Among such simulation-informed approaches, latent space models have shown promise. For example, EnsembleVAE [53], trained on MD snapshots and crystal structures of the soluble protein K-Ras, generates Cα ensembles from sampled latent features with full-atom picture subsequently produced by RoseTTAFold [4]. Similarly, idpSAM [26], trained on extensive simulations of intrinsically disordered regions (IDRs) with an implicit solvent model, produces Cα trace ensembles that can then be converted to all-atom representations. Such approaches demonstrate the power of leveraging simulation data within generative frameworks, though their application has often focused on particular protein classes (e.g., soluble proteins, IDRs) or involved multi-stage processes for generating final all-atom structures. Further advancements in learning ensembles directly from MD simulations encompass a range of techniques. These include methods like ConfDiff [27] (force-guided diffusion), P2DFlow [54] (SE(3) flow matching), and MD-Gen [55] (continuous trajectories). Larger-scale models such as BioEmu [22] and Distributional Graphormer (DiG) [23] aim to learn from vast MD datasets or equilibrium distributions, while experimentally-guided approaches like EGDiff [28] integrate diverse data types. Collectively, these MD-informed methods signify substantial progress in modeling protein dynamics. However, a persistent challenge, particularly when these methods are applied generally or to very large datasets, is the consistent generation of high-resolution, all-atom ensembles that fully capture system-specific details. This is especially true for intricate side-chain rearrangements within native environments, such as lipid membranes for G protein-coupled receptors (GPCRs), which are essential for elucidating their functional transitions [29]. The development of generative models that can directly learn and sample such specific, all-atom conformational landscapes from relevant MD data, particularly for complex targets like GPCRs [44], thus remains an important frontier.
@@ 3.1 Input Representation and Preprocessing
Each MD snapshot $t$ is represented by a graph $G^{(t)}=(V,E^{(t)})$, where the node set $V$ comprises the $N$ heavy atoms and the node features are their 3D coordinates $X^{(t)}\in\mathbb{R}^{N\times 3}$. For each frame, the edge index $E^{(t)}$ is built on the fly by applying a $k$-Nearest Neighbors search to the aligned coordinates $X^{(t)}$ with $k=4$, see Extended Technical Appendix I. The node positions serve both as input features and regression targets. Prior to graph construction, the raw MD coordinates $\tilde{X}^{(t)}$ are rigid-body aligned to the first frame using the Kabsch algorithm [56] to remove global rotation and translation.
@@ 3.2 Latent learning of conformations
Multi-hop encoding: For each MD frame $t$, the encoder $\Theta$ maps the Kabsch‑aligned heavy‑atom coordinates $X^{(t)}$ and their $k$‑NN graph $E^{(t)}$ to latent embeddings $Z^{(t)}\!\in\!\mathbb{R}^{N\times d_{z}}$, one $d_{z}$‑dimensional vector per heavy atom. We implement $\Theta$ as a four‑layer ChebNet [46] whose layers perform spectral graph convolutions with Chebyshev polynomials of order $K=4$:
@@ 3.4 Latent Diffusion Model for Generation
A Denoising Diffusion Probabilistic Model (DDPM) [47] operates on the pooled latent embeddings $\mathbf{h}_{0}$ derived from the chosen decoder pooling strategy. The model is trained to predict the noise $\epsilon$ that was added during a forward diffusion process. This training minimizes the standard DDPM loss function, $\mathcal{L}_{\text{diffusion}}$ (Eq. 2). New latent representations, $\mathbf{h}_{0}^{\text{gen}}$, are then sampled by iteratively applying the learned denoising network in a reverse diffusion process.
@@ 4.1 Experimental Setup
Implementation Details. All models were implemented in PyTorch [58] and trained using the Adam optimizer [59]. The overall three-phase training and generation workflow (encoder pre-training, decoder training, and diffusion model training) is detailed in Appendix A (Algorithm 1).
@@ 4.2 Results and Analysis
Furthermore, BioEmu [22], a general MD model, struggled to capture the D2R-specific conformational landscape: its generated A100 distributions (mean $-17.19\pm 7.45$) markedly diverged from our D2R-MD reference (mean $\approx-47.5\pm 2.26$), highlighting limitations of such generalist approaches for specific membrane protein ensembles.
@@ 5 Conclusion and Future Work
Future work includes targeted enhancements. For Blind and Sequential pooling, exploring larger pooled latent dimensions ($d_{p}$) is immediate, potentially improving detail capture but needing more extensive, diverse training data (e.g., from multiple related protein systems). For Residue pooling, with its promising high-dimensional effective latent space ($N_{\text{res}}\times d_{p}$), better denoisers are key to overcome multi-epoch sampling limits. This might use complex MLP/convolutional architectures, attention for its structured latent vector, or alternative generative models (score-based, flow-matching). For Sequential pooling, improving denoiser coverage of the MD-derived latent space (Fig.4b, backbone/sidechain overlap) could resolve side-chain misplacements and boost performance. Broader initiatives will enhance physical realism and architectural sophistication by incorporating lightweight energy surrogates, physics-guided diffusion schemes [60, 61, 28], and advanced architectures like attention-based fusion mechanisms [62, 63, 64], SE(3)-equivariant GNNs/transformers, or flow-matching generators [48, 65]. A long-term goal is generalizing LD-FPG by training on data from multiple related proteins (e.g., Class A GPCRs in various states) for foundational models akin to specialized “LLMs for protein dynamics.” Efforts will also cover tailoring pooling for applications, scaling to larger systems [18], and rigorous benchmarking.

### v4

@@ Abstract
Generating diverse, all-atom conformational ensembles of dynamic proteins such as G-protein-coupled receptors (GPCRs) is critical for understanding their function, yet most generative models simplify atomic detail or ignore conformational diversity altogether. We present latent diffusion for full protein generation (LD-FPG), a framework that constructs complete all-atom protein structures, including every side-chain heavy atom, directly from molecular dynamics (MD) trajectories. LD-FPG employs a Chebyshev graph neural network (ChebNet) to obtain low-dimensional latent embeddings of protein conformations, which are processed using three pooling strategies: blind, sequential and residue-based. A diffusion model trained on these latent representations generates new samples that a decoder, optionally regularized by dihedral-angle losses, maps back to Cartesian coordinates. Using D2R-MD, a $2\,\mu\text{s}$ MD trajectory (12 000 frames) of the human dopamine D22 receptor in a membrane environment, the sequential and residue-based pooling strategy reproduces the reference ensemble with high structural fidelity (all-atom lDDT $\sim 0.7$; $C\alpha$-lDDT $\sim 0.8$) and recovers backbone and side-chain dihedral-angle distributions with a Jensen–Shannon divergence <0.03<0.03 compared to the MD data. LD-FPG thereby offers a practical route to system-specific, all-atom ensemble generation for large proteins, providing a promising tool for structure-based therapeutic design on complex, dynamic targets. The D2R-MD dataset and our implementation are freely available to facilitate further research.
@@ 1 Introduction
Proteins function as dynamic molecular machines whose biological activities critically depend on transitioning between distinct conformational states (Karplus and Kuriyan, 2005; Henzler-Wildman and Kern, 2007). Landmark artificial intelligence methods like AlphaFold2 (Jumper et al., 2021) and others (Baek et al., 2021; Lin et al., 2023; Abramson et al., 2024; Wohlwend et al., 2024) have advanced structure prediction but predominantly predict single static conformations, limiting their utility for systems with conformational heterogeneity. Accurate modeling of an ensemble of accessible conformations is essential to elucidate protein function and guide therapeutic design (Boehr et al., 2009; Teague, 2003; Chen et al., 2020, 2025). Crucially, these ensembles must explicitly represent all atomic details, particularly side chains, whose subtle conformational rearrangements often govern molecular recognition and catalytic mechanisms Najmanovich et al. (2000).
@@
Despite rapid advancements, existing generative models frequently fall short in capturing the detailed dynamics of side-chain movements specific to particular proteins (Rudden et al., 2022). Many powerful methodologies have focused either on de novo backbone designs (Watson et al., 2023; Wu et al., 2024; Geffner et al., 2025) or on generating static all-atom structures (Chu et al., 2024; Ingraham et al., 2023), but neither produce comprehensive conformational ensembles. Ensemble-generating approaches—via perturbations of static predictors (Vani et al., 2023; Lu et al., 2023), flow-matched variants (Jing et al., 2024a), or general MD-trained generators (Lewis et al., 2024; Zheng et al., 2024)—typically operate at the backbone or coarse-grained level. Consequently, they fail to capture the intricate all-atom rearrangements, particularly involving side chains, that are critical for function. Although promising latent space (Fu et al., 2024; Lu et al., 2024; Janson and Feig, 2024) and physics-informed models (Wang et al., 2024; Liu et al., 2024) have emerged, their capability to generate high-resolution all-atom ensembles from MD data reflecting functional transitions remains unproven, highlighting a significant unmet need for specialized generative frameworks.
@@
G protein-coupled receptors (GPCRs) provide a compelling example of dynamic systems wherein precise all-atom modeling is indispensable (Latorraca et al., 2017a). This large family of transmembrane receptors, comprising over 800 human members, is responsible for mediating most known transmembrane signal transduction (Zhang et al., 2024; Casadó and Casadó-Anguera, 2023) and is targeted by approximately 50% of all marketed drugs (Yang et al., 2021; Wishart et al., 2018). GPCR signaling involves conformational transitions among multiple states, frequently induced by ligand binding (Hilger et al., 2018), and occurs through intricate allosteric mechanisms where specific side-chain interactions are crucial (Monod et al., 1965; Latorraca et al., 2017b). Capturing such dynamic, atomically detailed landscapes is vital to understanding receptor signaling specificity (Jefferson et al., 2023), biased agonism (Kenakin, 2019), and the design of drugs targeting unique allosteric sites (Goupil et al., 2012; Jeffrey Conn et al., 2009; Do et al., 2023), yet studying these events computationally remains formidable (Kapla et al., 2021; He et al., 2022; Rodr´ıguez-Espigares et al., 2020). Current predictive methods for GPCR dynamics (López-Correa et al., 2024) do not generate the comprehensive all-atom conformational landscapes essential for mechanistic understanding.
@@
To address this critical requirement, we introduce Latent Diffusion for Full Protein Generation (LD-FPG), a generative framework designed to learn and generate diverse, all-atom conformational ensembles from existing MD simulation data of a target protein, explicitly including side-chain details. Rather than simulating new trajectories, LD-FPG leverages extensive MD datasets. Our approach employs a Chebyshev Spectral Graph Convolutional Network (ChebNet) (Defferrard et al., 2016) to encode all-atom MD snapshots into a compact latent representation. A Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) is then trained to explore this learned latent manifold, and latent representations are decoded back into full all-atom Cartesian coordinates. We demonstrate our framework on extensive MD simulations of the human dopamine D2 receptor (D2R), systematically evaluating distinct decoder pooling strategies. Our primary contributions in this paper are:
@@ 2 Related Work
Generative Models for Protein Design and Static Structure Prediction. Deep generative models have made significant strides in protein science, employing techniques such as diffusion (Ho et al., 2020), flow-matching (Lipman et al., 2022; Huguet et al., 2024), and learned latent spaces Fu et al. (2024); Eguchi et al. (2022) to tackle complex structural tasks. For de novo backbone design, various methods, often inspired by the success of predictors like AlphaFold2 (Jumper et al., 2021) and diffusion-based approaches like RFdiffusion (Watson et al., 2023) have emerged. Notable examples include FoldingDiff (Wu et al., 2024), alongside flow-matching models such as FoldFlow2 (Huguet et al., 2024). Latent space strategies have also been pivotal in this domain; for instance, LatentDiff (Fu et al., 2024) generates novel protein backbones using an equivariant diffusion model within a condensed latent space, while Ig-VAE (Eguchi et al., 2022) utilizes a variational autoencoder for class-specific backbone generation (e.g., for immunoglobulins). While powerful for creating new folds (e.g., Proteina (Geffner et al., 2025)) or specific components, these methods typically target soluble proteins, often generate only backbone coordinates, and are not primarily designed for sampling multiple conformations of a specific, existing protein. The generation of complete static all-atom structures has also seen considerable progress. Models like Protpardelle (Chu et al., 2024) and Chroma (Ingraham et al., 2023) can produce full static structures from sequence information. Diffusion-based generative models such as AlphaFold3 (Abramson et al., 2024) and Boltz-1 (Wohlwend et al., 2024) also provide detailed single-state predictions of all-atom structures and complexes. Other approaches, like PLAID (Lu et al., 2024), integrate predictors with diffusion samplers. To complement backbone or static generation, methods including FlowPacker (Lee and Kim, 2025) and SidechainDiff (Liu et al., 2023) focus on side-chain packing or prediction. However, these tools predominantly yield single static structures. Furthermore, decoupling backbone and side-chain generation risks overlooking their critical interplay during the complex dynamic transitions (Latorraca et al., 2017a) relevant to the conformational landscapes our work aims to capture.
@@
Modeling Protein Conformational Diversity: From General Strategies to MD-Informed Approaches. Beyond single static structures, capturing a protein’s conformational diversity is crucial for understanding its function. Initial strategies to explore this diversity include learning from structural variations in experimental databases (e.g., Str2Str (Lu et al., 2023)) or perturbing static predictions to sample conformational space, as seen with AF2-RAVE (Vani et al., 2023) for GPCRs and AlphaFlow/ESMFlow (Jing et al., 2024a) more broadly. While these methods effectively broaden sampling, capturing system-specific, native-like dynamics often benefits from models trained directly on simulation data, which can provide a richer representation of a given protein’s accessible states. Among such simulation-informed approaches, latent space models have shown promise. For example, EnsembleVAE (Mansoor et al., 2024), trained on MD snapshots and crystal structures of the soluble protein K-Ras, generates Cα ensembles from sampled latent features with full-atom picture subsequently produced by RoseTTAFold (Baek et al., 2021). Similarly, idpSAM (Janson and Feig, 2024), trained on extensive simulations of intrinsically disordered regions (IDRs) with an implicit solvent model, produces Cα trace ensembles that can then be converted to all-atom representations. Such approaches demonstrate the power of leveraging simulation data within generative frameworks, though their application has often focused on particular protein classes (e.g., soluble proteins, IDRs) or involved multi-stage processes for generating final all-atom structures. Further advancements in learning ensembles directly from MD simulations encompass a range of techniques. These include methods like ConfDiff (Wang et al., 2024) (force-guided diffusion), P2DFlow (Jin et al., 2025) (SE(3) flow matching), and MD-Gen (Jing et al., 2024b) (continuous trajectories). Larger-scale models such as BioEmu (Lewis et al., 2024) and Distributional Graphormer (DiG) (Zheng et al., 2024) aim to learn from vast MD datasets or equilibrium distributions, while experimentally-guided approaches like EGDiff (Liu et al., 2024) integrate diverse data types. Collectively, these MD-informed methods signify substantial progress in modeling protein dynamics. However, a persistent challenge, particularly when these methods are applied generally or to very large datasets, is the consistent generation of high-resolution, all-atom ensembles that fully capture system-specific details. This is especially true for intricate side-chain rearrangements within native environments, such as lipid membranes for G protein-coupled receptors (GPCRs), which are essential for elucidating their functional transitions (Latorraca et al., 2017a). The development of generative models that can directly learn and sample such specific, all-atom conformational landscapes from relevant MD data, particularly for complex targets like GPCRs (Rodr´ıguez-Espigares et al., 2020), thus remains an important frontier.
@@ 3.1 Input Representation and Preprocessing
Each MD snapshot $t$ is represented by a graph $G^{(t)}=(V,E^{(t)})$, where the node set $V$ comprises the $N$ heavy atoms and the node features are their 3D coordinates $X^{(t)}\in\mathbb{R}^{N\times 3}$. For each frame, the edge index $E^{(t)}$ is built on the fly by applying a $k$-Nearest Neighbors search to the aligned coordinates $X^{(t)}$ with $k=4$, see Extended Technical Appendix I. The node positions serve both as input features and regression targets. Prior to graph construction, the raw MD coordinates $\tilde{X}^{(t)}$ are rigid-body aligned to the first frame using the Kabsch algorithm (Kabsch, 1976) to remove global rotation and translation.
@@ 3.2 Latent learning of conformations
Multi-hop encoding: For each MD frame $t$, the encoder $\Theta$ maps the Kabsch‑aligned heavy‑atom coordinates $X^{(t)}$ and their kk‑NN graph $E^{(t)}$ to latent embeddings $Z^{(t)}\!\in\!\mathbb{R}^{N\times d_{z}}$, one $d_{z}$‑dimensional vector per heavy atom. We implement $\Theta$ as a four‑layer ChebNet (Defferrard et al., 2016) whose layers perform spectral graph convolutions with Chebyshev polynomials of order $K=4$:
@@ 3.4 Latent Diffusion Model for Generation
A Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) operates on the pooled latent embeddings $\mathbf{h}_{0}$ derived from the chosen decoder pooling strategy. The model is trained to predict the noise $\epsilon$ that was added during a forward diffusion process. This training minimizes the standard DDPM loss function, $\mathcal{L}_{\text{diffusion}}$ (Eq. 2). New latent representations, $\mathbf{h}_{0}^{\text{gen}}$, are then sampled by iteratively applying the learned denoising network in a reverse diffusion process.
@@ 4.1 Experimental Setup
Implementation Details. All models were implemented in PyTorch (Paszke, 2019) and trained using the Adam optimizer (Kingma and Ba, 2014). The overall three-phase training and generation workflow (encoder pre-training, decoder training, and diffusion model training) is detailed in Appendix A (Algorithm 1).
@@ 4.2 Results and Analysis
Furthermore, BioEmu (Lewis et al., 2024), a general MD model, struggled to capture the D2R-specific conformational landscape: its generated A100 distributions (mean $17.19\pm 7.45$) markedly diverged from our D2R-MD reference (mean $\approx-47.5\pm 2.26$), highlighting limitations of such generalist approaches for specific membrane protein ensembles.
@@ 5 Conclusion and Future Work
Future work includes targeted enhancements. For Blind and Sequential pooling, exploring larger pooled latent dimensions ($d_{p}$) is immediate, potentially improving detail capture but needing more extensive, diverse training data (e.g., from multiple related protein systems). For Residue pooling, with its promising high-dimensional effective latent space ($N_{\text{res}}\times d_{p}$), better denoisers are key to overcome multi-epoch sampling limits. This might use complex MLP/convolutional architectures, attention for its structured latent vector, or alternative generative models (score-based, flow-matching). For Sequential pooling, improving denoiser coverage of the MD-derived latent space (Fig.4b, backbone/sidechain overlap) could resolve side-chain misplacements and boost performance. Broader initiatives will enhance physical realism and architectural sophistication by incorporating lightweight energy surrogates, physics-guided diffusion schemes (Wang and Chen, 2025; Zhang et al., 2025; Liu et al., 2024), and advanced architectures like attention-based fusion mechanisms (Lombard et al., 2025; Fang et al., 2025; Joshi et al., 2025), SE(3)-equivariant GNNs/transformers, or flow-matching generators (Lipman et al., 2022; Zhan et al., 2024). A long-term goal is generalizing LD-FPG by training on data from multiple related proteins (e.g., Class A GPCRs in various states) for foundational models akin to specialized “LLMs for protein dynamics.” Efforts will also cover tailoring pooling for applications, scaling to larger systems (Ingraham et al., 2023), and rigorous benchmarking.